ANNA HAS 12 BILLS IN HER WALLET, SOME $5 AND SOME $10 THE TOTAL VALUE OF THE BILLS IS $100. HOW MANY OF EACH BILL DOES ANNA HAVE?

Answer by checkley71(8403)

x+y=12 x=12-y
5x+10y=100 substitute (12-y) for x in the second equation & solve for y
5(12-y)+10y=100
60-5y+10y=100
5y=100-60
5y=40
y=40/5
y=8 there are 8 10 dollar bills
x+8=12
x=12-8
x=4 thus 4 5 dollar bills